# Diffusion Models for Image Restoration and Enhancement: A Comprehensive Survey

### v2

@@ 5.2 Experimental Results
To demonstrate the superiority of different diffusion models, we provide an objective quality comparison for them on multiple tasks. Concretely, we select three commonly investigated IR tasks, including image super-resolution, image deblurring, and image inpainting. The evaluation metrics are composed of PSNR, SSIM (Wang et al., 2004), FID (Heusel et al., 2017), and LPIPS (Zhang et al., 2018a). To compare the computation cost and network complexity, we also measure the running time, parameters, and flops for diffusion model-based IR methods. The subjective comparisons of image deblurring, image inpainting, and real-world IR are shown in Fig. 12, Fig. 3, 4, 5, 6 and 7 of section D.6 of the Supplementary. The comparison of diffusion-based methods on image inpainting and severe composite degradation can be found in the section D.4, D.5 of the Supplementary.
Results on Image SR and Real Image Restoration/SR. The experimental settings are as: (i) For the comparison of supervised diffusion model (DM)-based super-resolution (SR) methods, we use full images from DIV2K(Agustsson and Timofte, 2017) and Urban100(Huang et al., 2015) as test datasets instead of their cropped versions. We retrain SR3 (Saharia et al., 2022b) and SRDiff (Li et al., 2022b) with DF2K (Agustsson and Timofte, 2017) dataset to keep fair comparison. Additionally, we introduce two widely recognized diffusion-based Real-world SR methods, DiffBIR (Lin et al., 2024) and SUPIR (Yu et al., 2024), which are used to check the generalization capability of diffusion-based real-world image SR methods on synthetic datasets with bicubic degradation. (ii) For the comparisons of zero-shot DM-based SR methods, we follow the experimental setups of existing works, i.e., DPS(Chung et al., 2023b) and DDNM(Wang et al., 2023g), to process ImageNet1K (Russakovsky et al., 2015a) and CelebA-HQ test (Karras et al., 2017b) for consistent comparison. (iii) For the comparison of real-world image restoration, we follow StableSR (Wang et al., 2023c) to select RealSR (Cai et al., 2019), DRealSR (Wei et al., 2020), and RealPhoto60 as test datasets with central cropping operation.
@@
For zero-shot-based diffusion IR models, the quantitative comparisons are shown in Table 3. We test six open-source diffusion models for 4x super-resolution on two datasets: ImageNet 1K (Russakovsky et al., 2015a) and CelebaHQ test (Liu et al., 2015). From the table, we can observe that DDRM (Kawar et al., 2022a) and DDNM (Wang et al., 2023g) performed well on various metrics, followed by ILVR (Choi et al., 2021). This is because DDRM and DDNM consider data consistency with the low-resolution image from a decomposition perspective, while ILVR only ensures low-frequency consistency. DPS (Chung et al., 2023b) and GDP (Fei et al., 2023) focus more on the perceptual quality of generated images. GDP (Fei et al., 2023) performs well in perceptual metrics on both datasets because it uses an empirical formula rather than Tweedie’s formula for posterior estimation, resulting in some performance improvement relative to DPS. GDP also exhibits the fastest image generation speed, with the reverse sampling steps of 20. The reason for that is it does not need the complex SVD decomposition and computation in DDRM, thereby accelerating image generation. On the other hand, DPS performs corrections after every sampling step (around 1000 steps), making it unable to utilize the DDIM-based sampling method for acceleration. Therefore, generating a single image with DPS takes approximately 141.2 seconds.
@@
By comparing GAN-based IR methods and Diffusion-based methods, we can find that: (i) Diffusion-based IR methods do not always outperform GAN-based IR methods for synthetic degradation. As shown in Table 2 and 3, SeD (Li et al., 2024a) consistently achieve optimal reference-based perceptual quality metrics, including LPIPS and FID, which reveals the potential of Gan-based IR methods on synthetic degradation since they also have the advantage on efficiency; (ii) GAN-based IR methods and Diffusion-based IR methods are susceptible to poor pixel-wise consistency in general (i.e., PSNR, and SSIM) since they are generation-based models. But, we can find some zero-shot DM-based methods in Table 3, such as DDRM (Kawar et al., 2022a) and  (Wang et al., 2023g) achieve great pixel-wise consistency. The reason might be that decomposition-based methods can accurately obtain the inverse matrix for bicubic downsampling. Moreover, the utilized diffusion model for zero-shot IR has been pre-trained on face datasets; (iii) Diffusion-based RealSR methods outperform GAN-based SR methods by a large margin across all NR-IQA metrics for real-world degradation, as shown in Table 4, This could be attributed to the fact that these diffusion-based methods are built upon pre-trained diffusion models (e.g., Stable Diffusion (Wang et al., 2023c)) trained on large-scale natural image datasets. These models inherently possess strong priors that closely align with the real-world image distribution, enabling the super-resolution models fine-tuned on them to generate more realistic restoration results. Nevertheless, due to the nature of diffusion models, where images are sampled progressively from a Gaussian noise distribution, the stochastic generation process is difficult to constrain at the pixel level. This leads to inferior performance on objective metrics such as PSNR and SSIM compared to GAN-based methods, which can better preserve pixel-wise fidelity.
@@
We also select seven state-of-the-art typical deterministic IR methods with various architectures for comparison, including transformer (Dosovitskiy et al., 2020), mamba (Guo et al., 2024b) and CNNs (Zhang et al., 2018b) in Tables 2 and 3. These methods achieve superior performance on objective metrics such as PSNR and SSIM compared to GAN-based and diffusion-based methods. However, this strict pixel-wise fidelity constraint comes at the cost of perceptual quality, leading to inferior performance on subjective evaluation metrics compared to GAN-based and diffusion-based approaches.
@@
Results on Image Deblurring. We also evaluate five zero-shot DM-based IR methods on the Gaussian deblurring task using the ImageNet 1K (Russakovsky et al., 2015a) and CelebaHQ test (Liu et al., 2015) datasets. The experimental results are shown in Table 8 of the Supplementary. We can find that DiffPIR (Zhu et al., 2023) and Dirac-DO (Fabian et al., 2023) achieve competitive performance on PSNR and SSIM, with an average improvement of 1.0 dB to 1.4 dB over DDRM (Kawar et al., 2022a) and DDNM (Wang et al., 2023g). Moreover, Dirac-PO (Fabian et al., 2023) and DiffPIR (Zhu et al., 2023) show superior performance on perceptual metrics. DPS (Chung et al., 2023b) performs well on perceptual metrics, including LIPS and FID, but has a long generation time for each image. DiffPIR utilizes a plug-and-play sampling method and merges the DDIM sampling strategy to ensure both the fidelity and realism of generated images while speeding up the sampling process. Diracdiffion includes perception-optimized (PO) and distortion-optimized (DO) models, and employs an incremental reconstruction and early stopping method to achieve a perception-distortion trade-off. As a result, both models perform exceptionally well on distortion and perception metrics. Among all the models used, DDRM has the shortest sampling time, averaging under 10 seconds per image, as it only uses 20 sampling steps. In terms of model parameters, all zero-shot DM-based IR methods use a pre-trained model with 552.8M parameters for the ImageNet dataset and 126M parameters for the CelebA dataset.
@@
RealESRGAN (Wang et al., 2021b) 25.69 0.762 0.271 60.37 0.448 0.548 28.64 0.805 0.282 54.27 0.451 0.490 50.95 0.599 0.488 1
CAL-GAN (Park et al., 2023) 26.38 0.754 0.272 56.07 0.423 0.505 28.91 0.799 0.289 50.00 0.423 0.460 51.52 0.646 0.522 1
SeD (Li et al., 2024a) 25.56 0.754 0.271 61.03 0.540 0.546 28.35 0.794 0.285 55.75 0.530 0.500 57.14 0.690 0.519 1
CDC (Wei et al., 2020) 27.81 0.787 0.329 42.71 0.299 0.321 32.76 0.877 0.291 40.27 0.356 0.283 29.52 0.369 0.268 1
SwinIR (Liu et al., 2021) 26.87 0.800 0.291 53.73 0.306 0.349 29.06 0.834 0.329 45.42 0.330 0.307 30.82 0.332 0.244 1
MambaIR (Guo et al., 2024b) 26.72 0.766 0.265 53.38 0.396 0.299 29.01 0.805 0.302 45.35 0.386 0.288 31.46 0.351 0.253 1
StableSR (Wang et al., 2023c) 24.70 0.707 0.300 65.88 0.623 0.428 28.15 0.752 0.328 58.51 0.636 0.387 62.24 0.651 0.361 200
DiffBIR (Lin et al., 2024) 25.07 0.640 0.395 64.85 0.638 0.615 26.53 0.630 0.495 61.19 0.634 0.592 64.45 0.689 0.418 50
PASD (Yang et al., 2024b) 24.29 0.663 0.343 68.69 0.659 0.649 27.00 0.708 0.393 61.81 0.677 0.585 63.28 0.712 0.477 20
DiffIR (Xia et al., 2023) 26.52 0.767 0.256 58.60 0.416 0.351 29.18 0.808 0.271 50.54 0.426 0.311 49.41 0.621 0.353 4
ResShift (Yue et al., 2023) 25.69 0.736 0.328 56.89 0.536 0.347 27.12 0.741 0.387 51.24 0.539 0.325 36.41 0.566 0.289 15
SUPIR (Yu et al., 2024) 23.47 0.537 0.518 67.06 0.663 0.641 23.76 0.563 0.659 65.56 0.592 0.581 67.04 0.746 0.515 50
@@ 6.1 Sampling Efficiency
It is noteworthy that sampling efficiency is one typical challenge for the diffusion model, where the few sampling steps will cause limited generation fidelity. This inherent problem of the diffusion model damages the training and inference speed of image restoration. As shown in Table 2, SR3 (Saharia et al., 2022b) needs to take about 50 seconds to restore one image with the size of $224\times 224$, which is largely slower than the existing IR methods. previous works on diffusion models have attempted to improve the sampling efficiency from four perspectives: 1) modeling the diffusion process with a non-Markov Chain, such as DDIM (Song et al., 2020a). 2) designing efficient ODE solvers, e.g., DPM-solver. 3) leveraging the knowledge distillation to reduce sampling steps (Zhou and Tulsiani, 2023; Salimans and Ho, 2022), and 4) introducing the cross-modality priors with condition mechanism (Abu-Hussein et al., 2022; Wang et al., 2023c; Liu et al., 2023b). Under the above progresses, the sampling steps for the diffusion model are largely reduced to $10\sim 20$ steps, which also serves for more fast image restoration. In particular, DDRM (Kawar et al., 2022a) reduces the inference speed to 8 seconds for one $224\times 224$ of the image with the sampling strategy of DDIM (Song et al., 2020a).
Despite that, the above strategies are not specific to image restoration tasks. Differently, considering the low-quality image in IR contains abundant structural and textual information, some works (Luo et al., 2023b; Welker et al., 2022; Luo et al., 2023c; Zhao et al., 2023a) achieve image restoration by sampling from the low-quality image instead of pure noise, which obviates the extra sampling steps in the original DDPM. Among them, PartDiff (Zhao et al., 2023a) utilizes an intermediate latent stage containing Low-resolution information as a proxy for the start of the inference process. Another work (Ma et al., 2023) tries to accelerate the sampling speed of SR models by exploring the optimal boundary conditions for solving diffusion ODEs.
@@ 6.2 Model Compression
Model size is also a significant factor impacting the computational cost, limiting the real-time application of diffusion model-based image restoration (IR), such as mobile devices. In particular, DDPM (Ho et al., 2020) and SR3 (Saharia et al., 2022b) have 113.7M and 155.3M parameters, respectively, substantially exceeding previous CNN-based (Zhang et al., 2018b; Jiang et al., 2021a; Li et al., 2020) or Transformer-based IR backbones (Liu et al., 2021, 2022b; Wang et al., 2022e; Li et al., 2022a). To mitigate this, diffusion model compression is a potential but under-explored research direction toward efficient IR. Model compression (Li et al., 2023e), with the goal of reducing computational cost while maintaining task performance, has achieved great breakthroughs from four perspectives: including 1) model pruning, with the aim of removing the unimportant parameters by estimating the importance score of each parameter, 2) model quantization targets for reducing the bit-depth of the floating-point parameters for storage or computing, 3) knowledge distillation is proposed to transfer the knowledge from the complex teacher model to the simple and efficient student model, and 4) low-rank decomposition is devoted to decomposing the parameter tensor into multiple low-rank tensors. On basis of these, some works have taken a step forward to investigate the model compression of the diffusion model. Kim et al.(Kim et al., 2023a) introduce the block-removed knowledge distillation for the diffusion model, which constructs the student model by removing some residual and attention blocks from the U-Net architecture. Fang et al.(Fang et al., 2023) pinpoint that not all diffusion steps contribute to the generation process, and then, exploit partial diffusion steps to estimate the important and unimportant weights for parameter pruning with Taylor expansion. Additionally, there are pioneering works (Shang et al., 2023b; Wang et al., 2023a) that take the model quantization for the diffusion model to accelerate the sampling process. Although such progress, few works explore how to design the model compression for diffusion model-based IR, which is expected to be developed for real-time application.

### v3

@@ 5.2 Experimental Results
To demonstrate the superiority of different diffusion models, we provide an objective quality comparison for them on multiple tasks. Concretely, we select three commonly investigated IR tasks, including image super-resolution, image deblurring, and image inpainting. The evaluation metrics are composed of PSNR, SSIM (SSIM, ), FID (FID, ), and LPIPS (LPIPS, ). To compare the computation cost and network complexity, we also measure the running time, parameters, and flops for diffusion model-based IR methods. The subjective comparisons of image deblurring, image inpainting, and real-world IR are shown in Fig. 12, Fig. 3, 4, 5, 6 and 7 of section D.6 of the Supplementary. The comparison of diffusion-based methods on image inpainting and severe composite degradation can be found in the section D.4, D.5 of the Supplementary.
Results on Image SR and Real Image Restoration/SR. The experimental settings are as: (i) For the comparison of supervised diffusion model (DM)-based super-resolution (SR) methods, we use full images from DIV2K(DIV2K, ) and Urban100(huang2015single-urban100, ) as test datasets instead of their cropped versions. We retrain SR3 (SR3, ) and SRDiff (li2022srdiff, ) with DF2K (DIV2K, ) dataset to keep fair comparison. Additionally, we introduce two widely recognized diffusion-based Real-world SR methods, DiffBIR (lin2023diffbir, ) and SUPIR (yu2024scaling-SUPIR, ), which are used to check the generalization capability of diffusion-based real-world image SR methods on synthetic datasets with bicubic degradation. (ii) For the comparisons of zero-shot DM-based SR methods, we follow the experimental setups of existing works, i.e., DPS(dps, ) and DDNM(ddnm, ), to process ImageNet1K (imagenet1k, ) and CelebA-HQ test (Celeba-hq, ) for consistent comparison. (iii) For the comparison of real-world image restoration, we follow StableSR (wang2023exploiting-stable-sr, ) to select RealSR (cai2019toward-realsr, ), DRealSR (wei2020component-drealsr, ), and RealPhoto60 as test datasets with central cropping operation.
@@
For zero-shot-based diffusion IR models, the quantitative comparisons are shown in Table 3. We test six open-source diffusion models for 4x super-resolution on two datasets: ImageNet 1K (imagenet1k, ) and CelebaHQ test (liu2015deep-celeba, ). From the table, we can observe that DDRM (DDRMDRM, ) and DDNM (ddnm, ) performed well on various metrics, followed by ILVR (choi2021ilvr, ). This is because DDRM and DDNM consider data consistency with the low-resolution image from a decomposition perspective, while ILVR only ensures low-frequency consistency. DPS (dps, ) and GDP (GDP, ) focus more on the perceptual quality of generated images. GDP (GDP, ) performs well in perceptual metrics on both datasets because it uses an empirical formula rather than Tweedie’s formula for posterior estimation, resulting in some performance improvement relative to DPS. GDP also exhibits the fastest image generation speed, with the reverse sampling steps of 20. The reason for that is it does not need the complex SVD decomposition and computation in DDRM, thereby accelerating image generation. On the other hand, DPS performs corrections after every sampling step (around 1000 steps), making it unable to utilize the DDIM-based sampling method for acceleration. Therefore, generating a single image with DPS takes approximately 141.2 seconds.
@@
By comparing GAN-based IR methods and Diffusion-based methods, we can find that: (i) Diffusion-based IR methods do not always outperform GAN-based IR methods for synthetic degradation. As shown in Table 2 and 3, SeD (li2024sed, ) consistently achieve optimal reference-based perceptual quality metrics, including LPIPS and FID, which reveals the potential of Gan-based IR methods on synthetic degradation since they also have the advantage on efficiency; (ii) GAN-based IR methods and Diffusion-based IR methods are susceptible to poor pixel-wise consistency in general (i.e., PSNR, and SSIM) since they are generation-based models. But, we can find some zero-shot DM-based methods in Table 3, such as DDRM (DDRMDRM, ) and  (ddnm, ) achieve great pixel-wise consistency. The reason might be that decomposition-based methods can accurately obtain the inverse matrix for bicubic downsampling. Moreover, the utilized diffusion model for zero-shot IR has been pre-trained on face datasets; (iii) Diffusion-based RealSR methods outperform GAN-based SR methods by a large margin across all NR-IQA metrics for real-world degradation, as shown in Table 4, This could be attributed to the fact that these diffusion-based methods are built upon pre-trained diffusion models (e.g., Stable Diffusion (wang2023exploiting-stable-sr, )) trained on large-scale natural image datasets. These models inherently possess strong priors that closely align with the real-world image distribution, enabling the super-resolution models fine-tuned on them to generate more realistic restoration results. Nevertheless, due to the nature of diffusion models, where images are sampled progressively from a Gaussian noise distribution, the stochastic generation process is difficult to constrain at the pixel level. This leads to inferior performance on objective metrics such as PSNR and SSIM compared to GAN-based methods, which can better preserve pixel-wise fidelity.
@@
We also select seven state-of-the-art typical deterministic IR methods with various architectures for comparison, including transformer (dosovitskiy2020image-vit, ), mamba (guo2024mambair, ) and CNNs (zhang2018image-RCAN, ) in Tables 2 and 3. These methods achieve superior performance on objective metrics such as PSNR and SSIM compared to GAN-based and diffusion-based methods. However, this strict pixel-wise fidelity constraint comes at the cost of perceptual quality, leading to inferior performance on subjective evaluation metrics compared to GAN-based and diffusion-based approaches.
@@
Results on Image Deblurring. We also evaluate five zero-shot DM-based IR methods on the Gaussian deblurring task using the ImageNet 1K (imagenet1k, ) and CelebaHQ test (liu2015deep-celeba, ) datasets. The experimental results are shown in Table 8 of the Supplementary. We can find that DiffPIR (zhu2023denoising-diffpir, ) and Dirac-DO (fabian2023diracdiffusion, ) achieve competitive performance on PSNR and SSIM, with an average improvement of 1.0 dB to 1.4 dB over DDRM (DDRMDRM, ) and DDNM (ddnm, ). Moreover, Dirac-PO (fabian2023diracdiffusion, ) and DiffPIR (zhu2023denoising-diffpir, ) show superior performance on perceptual metrics. DPS (dps, ) performs well on perceptual metrics, including LIPS and FID, but has a long generation time for each image. DiffPIR utilizes a plug-and-play sampling method and merges the DDIM sampling strategy to ensure both the fidelity and realism of generated images while speeding up the sampling process. Diracdiffion includes perception-optimized (PO) and distortion-optimized (DO) models, and employs an incremental reconstruction and early stopping method to achieve a perception-distortion trade-off. As a result, both models perform exceptionally well on distortion and perception metrics. Among all the models used, DDRM has the shortest sampling time, averaging under 10 seconds per image, as it only uses 20 sampling steps. In terms of model parameters, all zero-shot DM-based IR methods use a pre-trained model with 552.8M parameters for the ImageNet dataset and 126M parameters for the CelebA dataset.
@@
RealESRGAN (wang2021realesrgan, ) 25.69 0.762 0.271 60.37 0.448 0.548 28.64 0.805 0.282 54.27 0.451 0.490 50.95 0.599 0.488 1
CAL-GAN (CAL-GAN, ) 26.38 0.754 0.272 56.07 0.423 0.505 28.91 0.799 0.289 50.00 0.423 0.460 51.52 0.646 0.522 1
SeD (li2024sed, ) 25.56 0.754 0.271 61.03 0.540 0.546 28.35 0.794 0.285 55.75 0.530 0.500 57.14 0.690 0.519 1
CDC (wei2020component-drealsr, ) 27.81 0.787 0.329 42.71 0.299 0.321 32.76 0.877 0.291 40.27 0.356 0.283 29.52 0.369 0.268 1
SwinIR (swinIR-transformer, ) 26.87 0.800 0.291 53.73 0.306 0.349 29.06 0.834 0.329 45.42 0.330 0.307 30.82 0.332 0.244 1
MambaIR (guo2024mambair, ) 26.72 0.766 0.265 53.38 0.396 0.299 29.01 0.805 0.302 45.35 0.386 0.288 31.46 0.351 0.253 1
StableSR (wang2023exploiting-stable-sr, ) 24.70 0.707 0.300 65.88 0.623 0.428 28.15 0.752 0.328 58.51 0.636 0.387 62.24 0.651 0.361 200
DiffBIR (lin2023diffbir, ) 25.07 0.640 0.395 64.85 0.638 0.615 26.53 0.630 0.495 61.19 0.634 0.592 64.45 0.689 0.418 50
PASD (yang2023pixel-PASD, ) 24.29 0.663 0.343 68.69 0.659 0.649 27.00 0.708 0.393 61.81 0.677 0.585 63.28 0.712 0.477 20
DiffIR (xia2023diffir, ) 26.52 0.767 0.256 58.60 0.416 0.351 29.18 0.808 0.271 50.54 0.426 0.311 49.41 0.621 0.353 4
ResShift (yue2023resshift, ) 25.69 0.736 0.328 56.89 0.536 0.347 27.12 0.741 0.387 51.24 0.539 0.325 36.41 0.566 0.289 15
SUPIR (yu2024scaling-SUPIR, ) 23.47 0.537 0.518 67.06 0.663 0.641 23.76 0.563 0.659 65.56 0.592 0.581 67.04 0.746 0.515 50
@@ 6.1 Sampling Efficiency
It is noteworthy that sampling efficiency is one typical challenge for the diffusion model, where the few sampling steps will cause limited generation fidelity. This inherent problem of the diffusion model damages the training and inference speed of image restoration. As shown in Table 2, SR3 (SR3, ) needs to take about 50 seconds to restore one image with the size of $224\times 224$, which is largely slower than the existing IR methods. previous works on diffusion models have attempted to improve the sampling efficiency from four perspectives: 1) modeling the diffusion process with a non-Markov Chain, such as DDIM (DDIM, ). 2) designing efficient ODE solvers, e.g., DPM-solver. 3) leveraging the knowledge distillation to reduce sampling steps (zhou2023sparsefusion-distillation, ; salimans2022progressive-distillation, ), and 4) introducing the cross-modality priors with condition mechanism (abu2022adir, ; wang2023exploiting-stable-sr, ; liu2023improved, ). Under the above progresses, the sampling steps for the diffusion model are largely reduced to $10\sim 20$ steps, which also serves for more fast image restoration. In particular, DDRM (DDRMDRM, ) reduces the inference speed to 8 seconds for one $224\times 224$ of the image with the sampling strategy of DDIM (DDIM, ).
Despite that, the above strategies are not specific to image restoration tasks. Differently, considering the low-quality image in IR contains abundant structural and textual information, some works (IR-SDE, ; welker2022driftrec, ; luo2023refusion, ; zhao2023partdiff, ) achieve image restoration by sampling from the low-quality image instead of pure noise, which obviates the extra sampling steps in the original DDPM. Among them, PartDiff (zhao2023partdiff, ) utilizes an intermediate latent stage containing Low-resolution information as a proxy for the start of the inference process. Another work (ma2023solving-ode, ) tries to accelerate the sampling speed of SR models by exploring the optimal boundary conditions for solving diffusion ODEs.
@@ 6.2 Model Compression
Model size is also a significant factor impacting the computational cost, limiting the real-time application of diffusion model-based image restoration (IR), such as mobile devices. In particular, DDPM (DDPM, ) and SR3 (SR3, ) have 113.7M and 155.3M parameters, respectively, substantially exceeding previous CNN-based (zhang2018image-RCAN, ; jiang2021towards-jpeg-old4-cnn, ; li2020learning-lab1, ) or Transformer-based IR backbones (swinIR-transformer, ; swinv2-transformer, ; wang2022uformer-transformer, ; li2022hst-lab3, ). To mitigate this, diffusion model compression is a potential but under-explored research direction toward efficient IR. Model compression (li2023modelcompressionSurvey, ), with the goal of reducing computational cost while maintaining task performance, has achieved great breakthroughs from four perspectives: including 1) model pruning, with the aim of removing the unimportant parameters by estimating the importance score of each parameter, 2) model quantization targets for reducing the bit-depth of the floating-point parameters for storage or computing, 3) knowledge distillation is proposed to transfer the knowledge from the complex teacher model to the simple and efficient student model, and 4) low-rank decomposition is devoted to decomposing the parameter tensor into multiple low-rank tensors. On basis of these, some works have taken a step forward to investigate the model compression of the diffusion model. Kim et al.(kim2023architecturalKDDiffusion_MC, ) introduce the block-removed knowledge distillation for the diffusion model, which constructs the student model by removing some residual and attention blocks from the U-Net architecture. Fang et al.(fang2023structuralprunningDiffusion_MC, ) pinpoint that not all diffusion steps contribute to the generation process, and then, exploit partial diffusion steps to estimate the important and unimportant weights for parameter pruning with Taylor expansion. Additionally, there are pioneering works (shang2023postQuantizationMC, ; wang2023towardsQuantizationMC, ) that take the model quantization for the diffusion model to accelerate the sampling process. Although such progress, few works explore how to design the model compression for diffusion model-based IR, which is expected to be developed for real-time application.